Instrument Log 

 

GRAMS RCF   C1 - Radiometer Calibration Facility     Lamont, OK

 

Date

Technician

Comments

9/1/00

KT

At 1755 GMT powered off GRAMS instrument per mentor request. The computer, electronics, UPS, and the pump were all powered off until further notice. RTS at 1810 GMT.

8/29/00

MR

Power cycled the computer because it had shut down due to the heat. Power cycled at 1550 GMT. RTS at 1605.

8/24/00

JEM

James Miles informed me that the GRAMSRCF had not collected since 8/23/00 @1500 GMT with the error message "Problems occurred during the collection process". Opened the instrument and power cycled the computer. The GRAMSRCF started collecting again at 1500 GMT on 8/24/00.

7/20/00

CAM

GRAMS RCF had not collected files since 1702 GMT on 7/17. Removed the GRAMS enclosure covers and found that the UPS power switch was on, but none of the LED indicators were lit. The enclosure temperature was 79F and the UPS should be operating at this temperature. Cycled the power switch and the UPS then began to operate. Sealed the enclosure and RTS at 1930 GMT. The following collection was successful at 2004 GMT.

7/17/00

CAM

SDS comlog reports indicated that GRAMS RCF had not collected files since 1402 GMT on 7/15. Removed GRAMS RCF enclosure covers at 1430 GMT and found that the Athena temperature controller was in an error state and locked up. As a result, the air conditioner was not running and the enclosure temperature was 95F. This had caused the UPS and computer to shut down. Power cycled the temperature controller to remove the error condition and checked all controller settings. Replaced the GRAMS covers and RTS at 1500 GMT. Note that the system remained off line until the air conditioner was able to cool the enclosure below 85F. Data collection resumed status O.K. at 1602 GMT.

7/13/2000

CM/JM

The TDDR shadow band was not rotating. Removed GRAMS enclosure covers at 1430 GMT and found that the program was running only the band was not moving. Measured the output from the TDDR shadow ring controller board and found it to be 0 volts between pins A and B. Measured between pin A and chassis ground and found it to be 24V. There was also 24 volts measured between pins B and chassis ground. This is the same problem that occurred on 3/28/00 when the controller board was found to be defective. Removed the controller board from GRAMS C1 and installed in GRAMS RCF. The shadow band motor would also not operate with this good board which indicated that the motor was defective as well. Removed the TDDR shadow band assembly and the controller board to the lab for T/S. T/S found that the motor was indeed bad and the controller board had a defective 2222A transistor. Replaced the transistor on the controller board. Notified the Program Manager and will await a decision on whether or not a replacement motor will be purchased. Installed the cover plate in place of the TDDR and system RTS at 1730 GMT with only the FSBR and TSBR operational.

6/26/00

SS

At 1700 GMT, the pump was observed as not operating. Several attempts were made to reset the pump with no avail. The pump was removed and replaced with a new one. After the replacement and some pressure adjustments were complete, the pump was observed to be in proper working order. RTS 1815 GMT. Status was OK at this time.

4/18/2000

CAM

Ken Black and John Smith were on site 4/17 to 4/18 to install the new GRAMS vacuum pump system. Installation was completed and the system required testing before activation. Tested the new GRAMS27 software and found that it runs OK in normal data acquisition mode. Tested the pump system by activating the manual pump down sequence. The sequence passed through the first two steps of switching the vacuum gauge on and opening valve 1, but encountered a turbo pump fault as step three (switch on turbo controller) was activated. Troubleshooting found that the turbo controller is emitting a constant 5V fault signal which does not allow the pump to come on and stops the sequence. Ken Black will do further testing in the laboratory at Sandia. Disabled the new GRAMS software and pump system and GRAMS RCF RTS running the old GRAMS25 software at 1810 GMT.

4/17/2000

Mentor

Ken Black and John Smith on site to install new GRAMS vacuum pump system. System to be off line most of the day. Also created new GRAMS26 and GRAMS27 directories so Scott Faas can FTP the new GRAMS software to the system. GRAMS26 will correct an error in the calculation of sunrise and sunset during daylight savings time. GRAMS27 contains the time calculation fix plus implements software for the automatic control of the vacuum pump system.

4/3/2000

MR

Received the band motor control card back from Tooman today. Anthony Bucholtz was onsite and installed it in the unit. The shadow band started working. RTS at 1510.

3/31/2000

CAM

Daily inspection on 3/27/00 found that the GRAMS RCF TDDR shadow band was not rotating. T/S on 3/28 traced the problem back to the shadow band controller board in the radiometer electronics unit. Continued T/S on 3/31 found that there is 24V present on both the positive and negative pins on the controller board resulting in 0V between the two. Powered GRAMS RCF radiometer electronics off at 1524 GMT. Removed the controller board from GRAMS C1 unit and installed in GRAMS RCF. The shadow band works fine with the C1 controller board. The negative pin is grounded on the new board and there is 24V between the positive and negative pins. Checked the components on the original board and was unable to locate the source of failure. Site Ops do not have the schematics for the controller board. Tim Tooman was on site and took the controller board with him to deliver to Scott Faas for repair. GRAMS RCF RTS at 1700 GMT without the shadow band controller. The TSBR and FSBR radiometers should be unaffected by this.

3/28/2000

CAM

Daily inspection on 3/27/00 found that the GRAMS RCF TDDR shadow band was not rotating. T/S on 3/28/00 found that the motor was not receiving 24VDC. Measured the voltage across the motor terminals to be 0V. Measure from the positive terminal to chassis ground and found that it was 24V. It appears as though the ground is floating somewhere. Traced the power connections all the way back to the shadow band controller board in the radiometer electronics unit and found the same result. 24V is present between the positive lead and chassis ground, but no voltage between the positive and negative leads. If the negative lead is grounded to chassis anywhere in the circuit the shadow band motor will operate. Further T/S postponed until after maintenance travel this week. System is operational but the shadow band is not.

3/27/2000

CAM

While troubleshooting a shadow band problem, GRAMS RCF temperature controller was observed to have failed. The different parameter levels cannot be accessed to check or set the values. R/R Athena *temperature controller. Set the parameters on the new controller as per instrument manual. RTS at 2215 GMT. *Old - 9829-80596 *New - 9829-80597

2/23/2000

CAM

The pump was not working. T/S found that the power toggle switch inside the pump enclosure was turned off. Restored power and RTS at 1700 GMT.

2/18/00

CAM

Received the new IOTech motherboard and digital I/O expansion board. Set up and installed the new components and placed into service. Powered the instrument on at 2110 GMT. Monitored all channels and everything appears to be working normally. System RTS at 2115 GMT.

2/8/00

CAM

GRAMS C1 IOTech unit was removed on 2/7 and installed in GRAMS RCF for testing purposes. The testing process is complete and the IOTech unit needs to be returned to GRAMS C1. Shut GRAMS RCF system down at 1500 GMT and removed the IOTech. Returned the IOTech to GRAMS C1. Left GRAMS RCF powered off. GRAMS RCF will remain off line until the new IOTech motherboard and digital I/O expansion board are received.

2/7/2000

CAM

GRAMS RCF radiometers were reinstalled on 2/2 after being modified by mentor. At that time it was observed that the system program would not run. The IOTech Daqbook 100 A/D interface was suspected to be the problem. Removed the IOTech unit from GRAMS RCF on 2/7. Inspection of the IOTech revealed that the Daqbook 100 motherboard and one of the DBK21 digital I/O expansion boards had sustained major water damage. It appears that the GRAMS enclosure had leaked water during the time that the radiometers were removed for calibration. Instrument mentor requested that Site Ops remove the IOTech from the GRAMS OT and install it in GRAMS RCF to confirm operation of GRAMS RCF and the newly modified radiometers. (GRAMS OT radiometers are out for modifications at this time and the system is off line). Removed GRAMS OT IOTech and installed in GRAMS RCF and powered the system on at 2000 GMT. GRAMS RCF immediately began to operate normally. Mentor requested that GRAMS RCF be left to operate in this manner overnight. Ken Black will be on site 2/8 to reinstall the GRAMS OT radiometers, and at that time the GRAMS OT IOTech will be returned to its original location and GRAMS OT will be brought back on line. Initiated a PR for new motherboard and digital I/O board to replace the failed units.

2/2/2000

Ken Black

Ken Black reinstalled all 3 radiometers at 1800 GMT and restarted the system. The system's internal monitor has failed. Site Ops connected an external monitor at 1930 GMT and found that the program was not running. LabView was open, but not updating. Verified the I/O Tech Daqbook was powered on. Rebooted the computer and restarted the software with no change in status. Advised Ken of the situation. Ken is looking into it. System remains off line at this time.

7/2/99

Mentor/KB

The sensors were removed by mentor for calibration.

6/29/99

CAM

SDS comlog reports indicate that GRAMS RCF has been shutting down for a couple of hours during the hottest part of the afternoons. This first began on 6/26 at 2101 GMT. The system recovers each day when the temperature begins to cool back off. Removed GRAMS RCF enclosure cover at 1315 GMT. Verified the Athena Temperature Controller settings were correct for levels 2, 3, and 5. Removed the air conditioner inlet filter and found substantial blockage. Removed the air conditioner exhaust filter and found that it too required cleaning. Cleaned both filters with dish soap and water. Air dried the filters and replaced in the system. Replaced cover and RTS at 1345 GMT. Will continue to monitor status to see if this solves the problem. Note that the land owner began cutting wheat on Friday 6/25 just a few yards west of the system and continued to cut around the site over the weekend.

6/2/99

CAM

GRAMS RCF radiometer deck was out of level. The bubble was nearly halfway out of the inner circle to the west. Loosened the GRAMS enclosure leveling nuts and leveled so that the bubble is centered within the inner circle. Secured the leveling nuts and RTS at 1530 GMT.

5/14/99

JEM

Performed weekly PM. RTS @1506 GMT.

4/16/99

MW

The air pump failed. The internal vanes were broken resulting in no air flow. Replaced the pump at 1800 GMT.

4/7/99

CAM

TDDR is not currently installed at this time. The ventilator pump has failed. Waiting for replacement pump.

3/18/99

MK

The ventilator pump has seized up. No spares to replace with at this time.

3/17/99

CAM

The 2001 GMT collection failed due to an FTP error encountered during the collection process. Attempts to ping the GRAMS computer remotely were unsuccessful. Removed the GRAMS enclosure front cover at 2040 GMT. Found that the Athena Temperature controller set point had reset to 100F. As a result the enclosure temperature was 100F which in turn had shut the PC down. Checked all other controller level settings to be correct. Corrected the set point to 75F at 2043 GMT and observed the air conditioner turn on. Temporarily replaced the enclosure cover at 2045 GMT. Removed the cover at 2055 GMT and observed the computer had powered back on. Replaced and sealed the enclosure cover. The following SDS collection was successful at 2101 GMT. RTS status O.K.

3/17/99

CAM

GRAMS RCF has been down since 10/5/98 due to problems with the IOTech A/D interface. Scott Faas and Mike Ferrario removed the interface to Sandia Labs for testing on 10/14/98 and found that the motherboard had failed. Received a replacement board from IOTech on 3/1/99. Shipped the A/D interface to Scott Faas for setup and testing of the new board on 3/4/99. Received the interface back at CF on 3/16/99 ready to be installed. Opened the GRAMS RCF enclosure and installed the interface and made all necessary connections at 1650 GMT on 3/17/99. Powered the system on at 1810 GMT. Verified the TSBR and FSBR were operational and connected to the proper ports. Replaced the enclosure covers and RTS status O.K. at 1850 GMT. Note that the TDDR sensor and motor are still out for repair.

11/13/98

Mentor

Scott Faas removed the A/D interface and is waiting on instructions from mentor whether to have repaired or replace. System is off line.

10/14/98

SF/MF

Scott Faas and Mike Ferrario on site and report the following: Booster fan to increase cooling flow to air conditioner installed. Insulation below radiometer deck installed. Software problem reported on this unit investigated and found to be due to failure of the IOTech analog/digital interface. A/D interface removed and returned to Sandia for diagnosis and repair. Radiometer electronics and thermoelectric heater/cooler controller were turned off. The computer, heater, UPS, fans, and air conditioner were left on.

10/5/98

CAM

SDS comlog reports indicated that no files had been received from GRAMS RCF since 10/5/98. The error message 'Collection Failed - No Files to Collect' had occurred since that time. Checked GRAMS RCF status remotely and found that the GRAMS25 software was running but none of the channels were updating. All channels reported 0. The error light was also lit. Restarted GRAMS25 software at 1344 GMT. This cleared the error light but all channels remained zero and were not updating. Checked the relay and TTL supply voltages at instrument. All voltages checked OK. Rebooted computer several times with no change in status. SDS collections were successful at 1400, 1500, and 1600 after restarting the program, but have failed each time since then. Notified instrument mentor of the situation.

9/21/98

CAM

SDS comlog reports indicated that no files had been collected from GRAMS RCF since 0001 GMT on 9/21/98. SDS is not able to connect to GRAMS RCF. Removed GRAMS RCF cover at 1455 GMT. Found that the computer was locked up and would not respond to any commands. Power cycled the computer at 1505 GMT. Windows 95 loaded in Safe Mode. Rebooted computer again at 1508 GMT. Restarted GRAMS software and RTS status O.K. at 1510 GMT.

9/17/98

CAM

SDS comlog report indicated that no files had been received from GRAMS Cal since 2201 GMT on 9/16/98. The message 'Collection Failed - No Files to Collect' has been generated since that time. Remote status check using SMS found that the GRAMS program run/shutdown switch was in the shutdown position. Restarted program at 1302 GMT. RTS at 1310 GMT.

9/11/98

SS

Weekly maintenance revealed that the pump pressure was low. Found the out filter to be leaking air around the seal. Turned off the pump tightened filter and restored power to the pump. Pump pressure now reads 8 PSI. RTS at 1715 GMT.

 9/8/98

CAM

SDS comlog reports indicated that no files had been received from GRAMS RCF since 0501 GMT on 9/8/98. Remote status check using SMS software found that the GRAMS25.vi run/shutdown switch was in the shutdown position. Restarted the program and RTS at 1303 GMT. The program restarted OK however the FSBR and TSBR TTL's do not appear to be activated. Both TTL boxes are reporting 6.666E-6 and the values are not changing. Shut the program down and restarted again at 1439 GMT with the same result. The system is operational but the TTL's do not appear to be. T/S continues.

9/4/98

SS

The pump pressure reads 5.5. When attempting to adjust the pressure found that the out filter was loose and leaking. Attempted to tighten the filter and it broke off. Removed the threaded portion of the filter and replaced the filter with a spare. RTS at 1600 GMT.

9/3/98

CAM

SDS comlog reports indicate that no files have been collected from GRAMS RCF since 0101 GMT on 9/3/98. The error message 'No Files to Collect' is generated. Checked GRAMS RCF status remotely using SMS software and found that the run/shutdown switch was in the shutdown position. Moved the switch to the run position at 1305 GMT. Restarted GRAMS25.Vi program and RTS status O.K. at 1310 GMT.

8/31/98

CAM

SDS comlog reports indicated that no files had been received from GRAMS RCF since 0201 GMT on 8/31/98. Remote status check using SMS software found that the Run/Shutdown switch was in the Shutdown position. Returned the switch to the Run position at 1325 GMT. RTS status O.K. at 1330 GMT

8/28/98

CM/GS

Instrument mentor requested that the GRAMS RCF air conditioner be removed and sent in to local heat and air firm for service to check operation and refrigerant charge. Air conditioner was removed on 8/26/98 at 1442 GMT. Received the air conditioner back on 8/28/98. The service technician advised that they could find nothing wrong with the air conditioner and the refrigerant was fully charged. At mentors request the pressure switch was changed from 24 to 30. Installed GRAMS RCF air conditioner from 1430 to 1600 GMT. Powered the system up at 1600 GMT. The exhaust fan would not operate when the compressor kicked on. Powered the system back down at 1610 GMT to install GRAMS OT. Returned to GRAMS RCF at 1915 GMT. Removed the air conditioner and found that the squirrel cage exhaust fan had slipped off track and would not turn freely. Corrected the problem and reinstalled air conditioner at 2030 GMT. Restored power at 2040 GMT with exhaust fan operational. At that time the enclosure temperature was over 100F so the UPS had cut AC power until the enclosure cooled down. System not yet operational as of this time.

8/26/98

CAM/MW

GRAMS RCF has consistently shut down during the heat of the afternoon due to the air conditioner being unable to keep the enclosure below 85F. Instrument mentor requested that the GRAMS air conditioner be removed and taken in for service. The compressor will be checked out and the refrigerant charged if necessary. In addition the pressure switch will be changed from 24 to 30. Removed GRAMS RCF enclosure covers at 1440 GMT. Powered the system down at 1442 GMT. Removed the air conditioner to be sent in for service. Replaced the covers at 1505 GMT and taped plastic over the open hole where the air conditioner inlet was originally. System off line until air conditioner is returned.

8/10/98

CAM

SDS comlog indicated that no files had been received from GRAMS RCF since 0104 GMT on 8/10/98. Remote status check via SMS software found that the run/shutdown switch was in the shutdown position. Moved switch to the run position and restarted software at 1312 GMT. RTS status O.K. at 1315 GMT.

8/3/98

CAM

GRAMS RCF has been running in degraded mode since 6/18/98. The system has been shutting down on hot afternoons due to inaccurate high internal temperature readings. At that time it was discovered that the Athena Temperature Controller had failed and a replacement controller was ordered. Received replacement controller. Removed GRAMS RCF enclosure cover at 1415 GMT on 8/3/98. R/R Athena Temperature Controller at 1430 GMT. Set Levels 5, 3, and 2 parameters to the proper values. Replaced cover and RTS at 1445 GMT.

7/17/98

CAM

SDS comlog indicated that no files had been collected from GRAMS RCF since 2301 GMT on 7/16. Checked instrument status remotely via SMS and found that the run/shutdown switch was in the shutdown position. Moved the switch to the run position and RTS at 1312 GMT.

7/13/98

CAM

The Instrument Status Page indicated that no files had been collected from GRAMS RCF since 0001 GMT on 7/13/98. The error message 'Collection Failed - No Files to Collect' is generated. Checked instrument status remotely using SMS software. Found that the run/shutdown switch was in the shut down position. Moved the switch to the run position and RTS at 1330 GMT.

7/10/98

CAM

The GRAMS RCF unit has not transferred files to SDS since 0101 GMT on 7/10/98. The error message 'Collection Failed - No Files to Collect' is generated. Remote status check via SMS software revealed that the run/shutdown switch was in the shutdown position. Moved the switch to the run position at 1324 GMT. The data on the display still did not update. Attempted a remote reboot but the computer hung up. Removed the GRAMS RCF enclosure cover at 1420 GMT. Found that the computer had indeed locked up on reboot. Performed a hard boot at 1425 GMT. System RTS status O.K. at 1427 GMT.

6/26/98

SS

PM revealed that the pump reading was low. Inspection of pump assembly revealed that the pump plug on the out side was loose and leaking air. Tightened the plug and the air pressure return to 7. Adjusted pressure to 8.5.

6/26/98

CAM

The GRAMS OT Athena temperature controller was installed in GRAMS RCF on 6/25 as part of T/S to isolate cause of erroneous temperature readings. Removed GRAMS OT controller from GRAMS RCF at 1344 GMT on 6/26 and replaced with the original GRAMS RCF controller at 1415. Initiated a PR to purchase a new controller. GRAMS RCF will continue to run in degraded mode until a new controller is received. This will result in GRAMS RCF shutting down for a few hours each day in the heat of afternoon due to erroneous controller temperature reading.

6/25/98

CAM

The GRAMS RCF has been shutting down during mid afternoons with a high temperature alarm. The internal process temperature reports 20-25 degrees higher than the actual internal temperature as checked with a calibrated handheld. Initial troubleshooting points towards a failed Athena controller or possibly the temperature sensor. In an effort to further isolate the failed component, the Athena controller was removed from the GRAMS RCF at 1908 GMT and installed in GRAMS OT at 1930 GMT. The GRAMS OT Athena controller was then installed in the GRAMS RCF at 1938 GMT. Preliminary indications now are that the GRAMS OT is showing the same characteristics as GRAMS RCF was originally and GRAMS RCF is operational. GRAMS OT has now shut down due to high temperature alarm. At this time it appears that it was the controller that was causing the erroneous temperature readings. Will allow systems to stabilize and continue to monitor both with current set up for 24 hours and order replacement parts as necessary.

6/22/98

CAM

SDS comlog report revealed that GRAMS RCF had not successfully collected data since 6/21/98 at 0401 GMT. Checked instrument status at 1500 GMT and found that the enclosure internal temperature was 100F while the actual outdoor temperature was only 80F. The high internal temperature causes the UPS power to be cut off. The temperature controller indicated that the enclosure heater was on. Pressed the Parameter/Access key on the controller at which point the heater went off. Checked all parameters and found that they were correctly set. Returned the controller to menu level 0 and returned to run mode at 1530 GMT. The air conditioner remained on as it should. SDS comlog revealed that a file was collected at 1603 GMT. Will continue to monitor status.

6/18/98-6/19/98

 

 

CAM

Instrument mentor indicates that the GRAMS1 temperatures are reporting high. Opened the GRAMS1 case at 1910 GMT on 6/18. The controller display showed LOAD - No and all LED indicators were lit. According to the manual all LED indicators should light momentarily upon initial power up but then go out. It appeared as though the controller had lost power at some point and then hung up in this state when it powered back on. Changed the No to Yes and the controller then appeared to start working. Checked all of the menu level settings and found several of them were incorrect, including the sensing element had changed to type J. Set all of the parameters as they should be at 1930 GMT. Once the controller started working the power was cut as it should be due to the temperature being above the alarm setting. Checked the status 6/19 A.M. from Web page and found that it still had not come back on line. Found after reading the controller manual that level 5 configuration parameters need to be set first because it can effect other menu levels. Checked all parameters again at 1330 GMT on 6/19 and found two parameters in level two that were incorrect. C.Spr (cool spread) was set to 20 instead of 2 and C.HYS (cool hysteresis) was set to 2 instead of 1. When setting the parameters Thursday we started with level 2 and these two parameters did not even show up. Apparently after making changes to level 5 last it affected level 2 as well. Reset all parameters beginning with level 5 at 1600 GMT. At that time the internal temperature was 91 so the power was off. Replaced the enclosure cover to allow it to cool. At 2000 GMT the GRAMS had still not come back on line. Check of enclosure interior found that the temperature was 82F. It appears that the thermoelectric cooler is not sufficiently cooling the enclosure. The temperature needs to be down to 80F before the UPS power will be restored. For the time being the upper alarm limit was changed to 85F. Notified mentor of actions and T/S will continue. 

3/3/98

CAM

Instrument mentor requested Site Ops check the GRAMS RCF Low Voltage Power Supply for proper relay and TTL output power. Measured the relay output to be 11.8V. However, the TTL output measured 0V. Removed the Low Voltage Power Supply cover and found that the positive and negative lugs were shorted together inside. Loosened the retaining nuts and increased the distance between the positive and negative lugs. Tightened the nuts and covered the lugs with electrical tape to prevent them from shorting together again. Measured the TTL output at 4.4V and the relay output at 11.8V. RTS at 1905 GMT.

2/18/98

CAM/TMG

Instrument mentor advised that no data had been received from GRAMS RCF since 1/26/98. Removed the GRAMS RCF front and rear covers and found that the GRAMS25 Interface Program was running but there were no data values displayed. Further T/S found that the power strip that supplies the relay and TTL power was powered off. Powered the strip on and rebooted the computer at 1842 GMT. The data values returned to normal at around 1845 GMT. Mentor then verified that the unit was sending data. RTS status O.K.

1/26/98

SS/GS

No power to UPS power outlet. Mentor stated that the outlet was wired nonstandard to allow the UPS to be controlled by a temperature sensor. The temperature sensor shuts down power to the UPS below 50 degrees Fahrenheit. Opened the covers at 2050 GMT. Powered up the heater, fans, and UPS. Leveled unit. Powered up Radiometers at 2105. Powered up the computer at 2110 GMT. Replaced covers at 2115 GMT. RTS at 2117 GMT.

1/16/98

GS

The shadow band on the TDDR quit working and condensation kept gathering on the inside of the dome. Pulled the exterior covers off of the unit at 2030 GMT. Powered off the radiometers at 2034 GMT. Noticed that the GRAMS program had encountered an error and was not running on the monitor. The UPS was running in battery mode and could not be monitored. Decided to proceed on with the TDDR removal. Removed the TDDR and placed the blank cover plate over the hole and tightened in place. The TDDR drive chain was noticed broke and the drive motors shaft was bent. Next tried to find the cause of the UPS running on battery power. Checked the power cord and fuse. Both O.K. Performed a system check of the UPS and it passed, but it still ran with battery. Checked the receptacle power outage and it checked bad. The upper plug in on the receptacle is bad, but the lower plug in on the same receptacle is good. Double checked this by placing the UPS plug in into the lower plug in and the UPS ran O.K. and on normal AC power. Did not have time to R/R the receptacle so it was decided to completely power down the system and power off the main breakers to the unit in the RCF. Replaced the external covers at 2215 GMT. Unit down until further notice.

10/23/97

TMG

Changed network address from 198.124.98.66 to 198.124.99.54 Added SMS software for remote monitoring. Work performed on Grams-Cal from 1430 GMT to 1500 GMT resulted in data loss from rebooting the computer several times. RTS at 1500 GMT.